OpenStudy (anonymous):

What is the domain of y=2x squared -4

hartnn (hartnn):

\(y=2x^2-4\) is a polynomial and all polynomials have domain as all real numbers as x can take any rela value :)

hartnn (hartnn):

**real value.

OpenStudy (anonymous):

I havevno idea what your talking about..

OpenStudy (anonymous):

How do you find it

hartnn (hartnn):

you know the meaning of domain ?

hartnn (hartnn):

i am here to give you the idea :)

hartnn (hartnn):

see, domain is the set of values which the variable 'x' can take. so, if i consider f(x) = 1/x , then x CANNOT take the value 0, right ? because the denominator cannot be =0 . hence the domain of this f(x) = 1/x is all real numbers except 0. makes sense so far ?

OpenStudy (anonymous):

Yeah

hartnn (hartnn):

one more example! \(f(x)=\sqrt x\) here x cannot take negative values. square root of a negative number is not real. so the domain of this f(x) is x>= 0 because x needs to be positive (or can be 0) does this one makes sense too ?

OpenStudy (anonymous):

No I suck at math and I'm finna fail this test

hartnn (hartnn):

which part did u not get ?